OpenStudy (yinshy):

A large box contains buttons with pictures of animals. There are four different animals. A button is drawn and replaced. In 100 trials, 35 buttons with cats were drawn. What is the probability that the next button will be a different animal? A. The experimental probability is 35%. B. The experimental probability is 65%. C. The theoretical probability is 35%. D. The theoretical probability is 65%.

OpenStudy (yinshy):

@Anaise

OpenStudy (yinshy):

@AloneS

OpenStudy (yinshy):

@alivejeremy

OpenStudy (yinshy):

need help!

OpenStudy (yinshy):

@Aveline

OpenStudy (aaronandyson):

so the probability of drawing a cat is 35/100 hence the probability of it being a different animal would be 100-35/ 100 = 65/100 or 13/20 in percentage probability of cat is 35/100=35% so different animal would be 100%-35%=65%

OpenStudy (yinshy):

but is it experimental or theiroetical?

OpenStudy (yinshy):

@AaronAndyson

OpenStudy (alivejeremy):

One min let me post my question. :)

OpenStudy (yinshy):

ok

OpenStudy (mathmate):

Experimental probability bases on experimental data (hence could be different between experiments). Theoretical probability is based on theoretical calculations, and no experiment is required. So it does not vary. See if you can decide which one it is here.

OpenStudy (aaronandyson):

The theoretical probability is 65%. If we use the formulas and the theorems etc.

OpenStudy (alivejeremy):

The experimental probability is 65%.

OpenStudy (yinshy):

i say its B

OpenStudy (yinshy):

@Vocaloid